# SeedVR2: One-Step Video Restoration via Diffusion Adversarial Post-Training

### v1

@@ 5 Conclusion
Acknowledgement: This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2022-01-033[T]). We sincerely thank Zhibei Ma for data processing, Jiashi Li for hardware maintenance. We also give thanks to Feng Cheng for performance feedback, Yi Li, Fangyuan Kong and Rui Zhang for training suggestions.
Appendix A Limitations and Societal Impacts
To the best of our knowledge, SeedVR2 is the first one-step diffusion model towards video restoration (VR). Its effectiveness has also been verified with extensive experiments in the main paper. However, we further identify several limitations of current SeedVR2 in practice. We also discuss the potential societal impacts.
Limitations. While our one-step method significantly saves time during sampling, the causal video VAE requires over 4x more time to encode and decode a video compared to the naive VAE commonly used by existing methods [79, 97, 14, 74]. In addition, when dealing with a 720p video with 100 frames, the casual video VAE takes over 95% of the total time. How to effectively improve the efficiency the video VAE without significantly sacrificing the performance should be a valuable future work.
Besides the VAE efficiency, we notice that our method is sometimes not robust to heavy degradations and very large motions, and shares some failure cases with existing methods, e.g., fail to fully remove the degradation or simply generate unpleasing details. Moreover, due to the strong generation ability, SeedVR2 tends to overly generate details on inputs with very light degradations, e.g., 720p AIGC videos, leading to oversharpened results occasionally. Thus, we have to tune the model with careful hyperparameter settings. Improving the robustness of the model towards complex real-world degradations and ensuring a satisfactory lower bound of performance remains a challenge for future work.
@@ Appendix A Limitations and Societal Impacts
Societal Impacts. Our approach is likely to push forward the progress of restoration applications toward real-world image and video restoration. Specifically, our approach may inspire future work to develop fast restoration methods with strong performance. The release of our model weights and validation sets could further contribute to the restoration community in developing their own large restoration models. Of particular concern is the misconduct of applying our method to enhance illegal content, such as NSFW. To mitigate this risk, we plan to include the corresponding detection tool in our public code to restrict the use of our method.
Appendix B Parameter Size and Inference Speed
@@
We provide a detailed statistic regarding the number of parameters and inference time in Figure 1 of the main paper. We apply 50 sampling steps and keep other settings the same as the official repo for other baselines to maintain high-quality generation results of these methods. The results are listed in Table 4 for reference.
@@
Metrics VEnhancer UAV MGLD-VSR STAR SeedVR-7B Ours-7B Ours-7B
@@
Appendix C Additional Results
[…]
figure Qualitative comparisons on both real-world [6] and AIGC videos. It is noticeable that the GAN-based approach [95] generates blurry results due to limited generation ability. Previous multi-step diffusion-based VR [14, 97, 79, 74] either fail to restore the low-quality video with faithful details or tend to generate oversharpened details. Even with a single sampling step, our approach clearly excels over these methods with a large margin. (Zoom-in for best view).
We show additional comparisons in Figure C. For more image and video demos generated by our SeedVR2, please refer to the video demo in the project page: https://iceclear.github.io/projects/seedvr2/ for details.

### v2

@@ 5 Conclusion
6 Ethics Statement
@@
Our approach is likely to push forward the progress of restoration applications toward real-world image and video restoration. Specifically, our approach may inspire future work to develop fast restoration methods with strong performance. The release of our model weights and code could further contribute to the restoration community in developing their own large restoration models. Of particular concern is the misconduct of applying our method to enhance illegal content, such as NSFW. To mitigate this risk, we plan to include the corresponding detection tool in our public code to restrict the use of our method.
7 Reproducibility Statement
Our implementation is based on PyTorch 2.4.0 with CUDA 12.4. All the referenced software and models used in this paper are publicly available. Our code and model checkpoints will be released for reproducibility upon acceptance.
Acknowledgments
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2022-01-033[T]). We sincerely thank Zhibei Ma for data processing, Jiashi Li for hardware maintenance. We also thank Feng Cheng for performance feedback, Yi Li, Fangyuan Kong and Rui Zhang for training suggestions.
Appendix A Appendix
A.1 Use of Large Language Models
The large language models (LLMs), i.e., GPT-4o and Gemini 2.5 Pro, are solely used for polishing some paragraphs in this paper for clarity of expression and avoidance of minor grammar errors. They are not involved in any aspects related to the research contributions of this paper.
A.2 Additional Evaluations
Table 4: Additional full-reference metrics on synthetic benchmarks (SPMCS, UDM10, REDS30, YouHQ40). The best and second performances are marked in red and orange, respectively.
Datasets Metrics Bicubic RealViformer MGLD-VSR UAV VEnhancer STAR SeedVR-7B Ours-3B Ours-7B
SPMCS NIQE $\downarrow$ 9.105 3.431 3.315 3.272 4.328 5.659 3.671 3.862 3.668
MUSIQ $\uparrow$ 24.65 62.09 65.25 65.01 54.94 37.74 70.03 63.52 63.37
CLIP-IQA $\uparrow$ 0.3448 0.4239 0.4948 0.5074 0.3341 0.2346 0.4690 0.4736 0.4367
DOVER $\uparrow$ 0.9490 7.664 8.471 6.237 7.807 3.728 10.02 9.754 9.333
$E^{*}_{warp}$ $\uparrow$ 0.395 0.655 1.414 2.188 0.768 0.479 1.797 0.715 0.845
VMAF $\uparrow$ 5.24 20.43 34.47 27.19 16.96 14.13 43.05 39.96 39.86
UDM10 NIQE $\downarrow$ 8.625 3.922 3.814 3.494 4.883 5.273 4.025 4.545 4.518
MUSIQ $\uparrow$ 22.70 55.60 58.01 58.31 46.37 38.62 62.49 56.59 53.28
CLIP-IQA $\uparrow$ 0.3377 0.3972 0.4430 0.4583 0.3035 0.2338 0.4404 0.3695 0.3459
DOVER $\uparrow$ 1.592 7.259 7.717 9.238 8.087 5.374 10.63 9.652 8.907
$E^{*}_{warp}$ $\uparrow$ 0.506 0.630 1.412 1.558 0.771 0.775 1.953 0.777 0.830
VMAF $\uparrow$ 12.40 36.24 45.85 29.45 19.76 21.37 52.76 52.19 52.13
REDS30 NIQE $\downarrow$ 9.058 3.032 2.550 2.561 4.615 5.548 3.462 4.034 3.825
MUSIQ $\uparrow$ 19.43 58.60 62.28 56.39 37.95 30.08 57.80 50.90 50.79
CLIP-IQA $\uparrow$ 0.2538 0.3920 0.4442 0.3978 0.2445 0.2109 0.3019 0.2601 0.2622
DOVER $\uparrow$ 1.459 5.229 6.544 5.234 5.549 3.843 6.795 6.365 6.002
$E^{*}_{warp}$ $\uparrow$ 0.852 1.619 3.897 5.366 2.131 1.841 4.854 4.010 4.084
VMAF $\uparrow$ 9.21 28.32 38.05 28.16 12.05 18.88 43.28 41.73 39.93
YouHQ40 NIQE $\downarrow$ 9.241 3.172 3.255 3.000 4.161 5.470 3.288 3.722 3.378
MUSIQ $\uparrow$ 24.78 61.88 63.95 64.45 54.18 37.40 69.68 63.34 66.01
CLIP-IQA $\uparrow$ 0.3162 0.4376 0.5085 0.4710 0.3518 0.2466 0.5132 0.4649 0.4687
DOVER $\uparrow$ 1.276 9.483 10.50 9.957 11.44 3.817 13.42 13.34 12.98
$E^{*}_{warp}$ $\uparrow$ 0.606 0.996 2.057 3.249 1.284 0.785 3.210 1.376 1.725
VMAF $\uparrow$ 4.82 20.34 30.74 16.29 12.25 13.02 34.56 36.49 34.07
Additional Metrics. Due to the limited space, we provide additional metrics, i.e., NIQE (Mittal et al., 2012), MUSIQ (Ke et al., 2021), CLIP-IQA (Wang et al., 2023), DOVER (Wu et al., 2023), warping error (Zhou et al., 2024), and VMAF (Li et al., 2016) on the synthetic datasets (Yi et al., 2019, Tao et al., 2017, Nah et al., 2019, Zhou et al., 2024) for further evaluation across various baselines. As shown in Table 4, our proposed one-step approach consistently achieves comparable performance with other multi-step methods (50 steps) in terms of CLIP-IQA, MUSIQ on three of the four datasets, and DOVER, VMAF across all the four datasets. Note that the good performance of MGLD-VSR on REDS30 is expected since this method is trained on the training set of REDS (Nah et al., 2019) while other methods do not include such data for training. While our method does not show advantages in warping error, it is noticeable that the naive bicubic upsampling shows the best performance over all other existing baselines (i.e., 0.395, 0.506, 0.852, 0.606 on SPMCS, UDM10, REDS30, and YouHQ40, respectively). The reason is that the warping error measures the accuracy of the optical flow between the predicted frames and the ground-truth ones. Given the ill-posed nature of the restoration problem, the predictions may not perfectly align with the assigned ground-truth in synthetic datasets. Such a phenomenon can be more pronounced for models with stronger generative ability like ours. Thus, we argue that such a metric does not accurately reflect the temporal quality of our approach. Moreover, our method shows superior performance on other video metrics such as VMAF and DOVER, which both take temporal stability and temporal coherence into consideration. The user study in Table 2 and the demos provided in this supplementary material further demonstrate the high perceptual quality of our generated results.
Figure 6: We follow UAV (Zhou et al., 2024) to examine a row and track changes over time. Our method can achieve satisfactory temporal consistency with clear edges and patterns along the temporal dimension. (Zoom-in for best view).
Temporal Profile. We follow UAV (Zhou et al., 2024) to provide the visualization of the temporal profile in Figure 6. Compared with other baselines, our method can achieve satisfactory temporal consistency with clear edges and patterns along the temporal dimension.
Additional Visual Results. We further show additional comparisons in Figure 7. For more image and video demos generated by our SeedVR2, please refer to our project page: https://iceclear.github.io/projects/seedvr2/ for details.
Figure 7: Qualitative comparisons on both real-world (Chan et al., 2022a) and AIGC videos. It is noticeable that the GAN-based approach (Zhang and Yao, 2024) generates blurry results due to limited generation ability. Previous multi-step diffusion-based VR (He et al., 2024a, Zhou et al., 2024, Yang et al., 2024, Xie et al., 2025) either fail to restore the low-quality video with faithful details or tend to generate oversharpened details. Even with a single sampling step, our approach clearly excels over these methods with a large margin. (Zoom-in for best view).
Adversarial Training vs. Distillation. While distillation (Salimans and Ho, 2022) is a widely used strategy for diffusion acceleration, the effectiveness of distillation from a pre-trained teacher model in video restoration is still in doubt without further evidence. Without a teacher model, we observed that the model trained with our proposed adversarial approach presents the potential to surpass the initial model, which theoretically cannot be achieved with naive distillation from a teacher model.
Besides the inherent performance constraint imposed by the teacher model, the loss functions commonly used in distillation further limit the performance gains of the student model. In contrast, our method is based on adversarial training, which have been widely recognized as more effective for image restoration compared to non-adversarial methods, since SRGAN (Ledig et al., 2017). Compared to adversarial approaches, distillation from a teacher model resembles non-adversarial methods that rely on L1 or L2 losses, which tend to produce over-smoothed results with a performance upper bound from the teacher model, especially for sampling with very few steps. To further strengthen such a claim, we present a comparison between the distillation baseline without adversarial post-training and our proposed approach. Both of the baselines are trained for the same iterations for fair comparison. Both quantitative results in Table 5 and qualitative visualizations in Figure 8 on real datasets demonstrate the advantages of avoiding the constraint of a teacher model.
Figure 8: We present the visual comparison between progressive distillation (Salimans and Ho, 2022) and our proposed method based on adversarial training. The sharper edges and finer details indicate the potential of our proposed approach to avoid the constraint of a teacher model compared with the naive distillation. (Zoom-in for best view).
Table 5: Quantitative comparisons on real data between progressive distillation from a teacher model and ours.
Datasets Methods NIQE $\downarrow$ MUSIQ $\uparrow$ CLIP-IQA $\uparrow$ DOVER $\uparrow$
VideoLQ Prog. Distill. 5.365 45.57 0.230 6.609
Ours-3B 4.687 51.09 0.295 8.176
AIGC28 Prog. Distill. 4.857 58.85 0.416 13.11
Ours-3B 3.801 62.99 0.561 15.77
A.3 Parameter Size and Inference Speed
@@
We provide a detailed statistic regarding the number of parameters and inference time in Figure 1 of the main paper. We apply 50 sampling steps and keep other settings the same as the official repository for other baselines to maintain high-quality generation results of these methods. The results are listed in Table 6 for reference.
Figure 9: Comparison with DOVE-5B (Chen et al., 2025b) on real-world data. Our 3B model is capable of generating faithful details. (Zoom-in for best view).
@@
Metrics VEnhancer UAV MGLD-VSR STAR SeedVR-7B Ours-3B Ours-7B
@@
Table 7: Quantitative comparisons on real data between DOVE-5B and Ours-3B.
Datasets Methods NIQE $\downarrow$ MUSIQ $\uparrow$ CLIP-IQA $\uparrow$ DOVER $\uparrow$
VideoLQ DOVE-5B 5.079 50.91 0.341 8.389
Ours-3B 4.687 51.09 0.295 8.176
AIGC28 DOVE-5B 4.486 62.62 0.610 16.13
Ours-3B 3.801 62.99 0.561 15.77
A.4 Comparison with Concurrent Work
Before the submission, we further notice that there is a concurrent work (Chen et al., 2025b) targeting at one-step video restoration. Specifically, the proposed method, named as DOVE (Chen et al., 2025b) proposes a two-stage strategy where the VAE model is first adapted to achieve the restoration mapping between low-quality input and the ground truth, and the diffusion transformer model initialized from CogVideoX1.5 (Yang et al., 2025) is refined in the pixel domain with a combination of L1 loss, DISTS loss, and a frame difference loss. Our proposed approach considers an orthogonal path to achieve one-step acceleration with adversarial training and avoid training in the pixel domain for high efficiency under large-scale training. We make a comparison with DOVE-5B on real-world benchmarks in Table 7. Our approach with 3B parameters shows superior performance on the metrics, including NIQE and MUSIQ, but is slightly inferior according to CLIP-IQA and DOVER. It is noticeable that DOVE has 5B parameters, i.e., about 1.67x larger than our 3B model. We further provide visual comparisons in Figure 9. Our 3B model is capable of generating more faithful details compared with DOVE.
Figure 10: Quantitative comparison of speed-quality tradeoff (Zoom-in for best view).
Figure 11: Qualitative comparison with DLoRAL (Sun et al., 2025) on real-world videos. Video comparison can be found in the demo video in the supplementary materials (Zoom-in for best view).
A.5 Speed-quality Trade-off
We further provide metric performance of current multi-step diffusion-based restoration baselines under different sampling steps. As shown in Figure 10, with fewer sampling steps, the performance of most existing multi-step diffusion-based baselines (Yang et al., 2024, Zhou et al., 2024, He et al., 2024a, Xie et al., 2025) drops significantly in terms of perceptual metrics, including LPIPS, DISTS, MUSIQ, DOVER, and VMAF. Such speed-quality trade-off curves indicate that it is non-trivial to develop high-quality one-step diffusion-based methods for video restoration.
A.6 Comparison with DLoRAL
We additionally compare with another concurrent work DLoRAL (Sun et al., 2025), as shown in Figure 11. While DLoRAL focuses on parameter-efficient finetuning a pretrained diffusion prior (Rombach et al., 2022) via LoRA (Hu et al., 2022b), it mostly relies on the frozen generative prior, which is not specifically designed for restoration tasks. Our work alternatively explores the potential improvement on real-world restoration brought by large-scale training with sufficient compute, which is underexplored by previous methods. The improvement brought by our proposed approach is evident in Figure 11, with superior textures of the sofa, skin of the woman, and details of the building and boat. The temporal consistency is also better with less texture flicking, as shown in the demo video of the supplementary material.